OpenStudy (anonymous):

A 104 Ω resistor is joined in parallel with a 49 Ω resistor. What is the equivalent resistance? I tried adding the two and also adding 1/104 + 1/49 but I cant get the right answer

OpenStudy (anonymous):

ok so Req for parallel resistors = (1/r1 +1/r2)^-1 = (# of resistors/ resistance)^-1 = (resistance/#of resistors in parallel) = Req

OpenStudy (anonymous):

Req = (1/104 + 1/49)^-1 = Req (plug into your calc)

OpenStudy (anonymous):

((1 / 104) + (1 / 49))^(-1) = 33.3071895

OpenStudy (anonymous):

if they were a series you would just add them strait out Req = R1 + R2...

OpenStudy (anonymous):

its that the quantity is calculated for parallels to the -1 power, and you were missing that

OpenStudy (anonymous):

Your answer is 1/Req: \[\frac{1}{R_{eq}} = \frac{1}{R_1} + \frac{1}{R_2}\]

OpenStudy (anonymous):

thanks!